Lee’s mother was 3 times Lee’s age six years ago. Now Lee’s mother is 28 years older than
Lee. What is Lee’s age now?

Answer by fractalier(6550)

Call Mom's and Lee's age NOW, M and L.
Then we can write
M - 6 = 3(L - 6) and
M = L + 28
Substitute this into the first equation and get
L+28 - 6 = 3(L - 6)
L + 22 = 3L - 18
Add 18 and subtract L from both sides and get
L + 40 = 3L
40 = 2L
L = 20 years old NOW.